which of the expressions below are equal to


"4\u20221 + 7\u20224 + 10\u20229 +13\u202216..."


when summed to the "k" terms?


1)

"k"

"\u03a3" "(a^3 + 3a^2)"

"a = 1"


2)

"k"

"\u03a3" "(3a-1)(a^2)"

"a = 1"


3)

"k"

"\u03a3" "(3a^3 + a^2)"

"a = 1"



4)

"\ufeffk(k+1)(9k^2 + 13k +2)"

_________________________

"12"


5)


"k(k+1)(9k^2+2k+13)"

___________________________

4


1
Expert's answer
2021-06-07T18:40:37-0400

3)

"\displaystyle\sum_{a=1}^{k}(3a^3+a^2)=\displaystyle\sum_{a=1}^{k}((3a+1)a^2)"

"=4\cdot1+7\cdot4+10\cdot9+13\cdot16+...(3k+1)k^2"